OpenStudy (anonymous):

betsy bought a toaster over for $90. the price was 1/4 off the list price. what was the list price?

OpenStudy (anonymous):

Besty bought a toaster oven for 90 dollars. this price was 1/4 off the list price. what was the list price. List price = x 1/4 of list price = x(1/4)=x/4 List pirce - discount = Discounted price x-x/4=90 4x-x/4=90 3x/4=90 x=90*4/3 x=30*4 So all you have to do if figure out what is 30 * 4